# Galaxy mass profiles with convolutional neural networks

### v1

@@ 4 Neural Network
We use the python’s package tensorflow (Abadi et al., 2015) for constructing a CNN based on the one used in Expósito-Márquez et al. (2023).
@@ 4.1 Input
We construct the inputs using the x and y line-of-sight projected positions of stars in the sky, and their velocities along the line-of-sight , for each projection of each galaxy in the dataset. Additionally we use the projected half-light radius , the 98th percentile of the line-of-sight velocity , and the line-of-sight velocity dispersion of stars . We calculate these quantities for each projection using only the stars in the sub-sample corresponding to the projection, rather than all stars in the galaxy.
We choose to use two different 2-d arrays with dimensions 64x64 as inputs. The arrays represent samples of 2-d joint probability distribution functions (PDFs) of projected stellar data. We construct the PDFs using bivariate kernel density estimation (KDE), which is described in Appendix A. We find that these maps impact more positively the training of the model when we construct them using only stars within a radius of 0.5. The input arrays are described next.
Input 1: A 2-d PDF of the projected positions of stars in the sky, normalised by (The two dimensions of the array are {x/, y/}). We sample the PDF in a grid between -0.5 and 0.5 in both dimensions to create a 64x64 array.
Input 2: A 2-d PDF in the space {/, /} (where ). We sample the PDF in a grid between -0.5 and 0.5 in the / dimension and between -1 and 1 in the / dimension, obtaining a 64x64 array.
Input 3: Since literature mass estimators make use of line-of-sight velocity dispersion and projected half-light radius , we decide to also input these values to the final layers of the network. On top of that, we also include the value , as we have used it for scaling one dimension of the array in input 2. We structure this data in a one-dimensional array of size 3.
@@ 4.2 Architecture
The architecture of the CNN is shown in in Fig. 3. It is formed by two branches which take as inputs both arrays described in section 4.1. The first input (projected position map {x/, y/}) is randomly rotated and/or flipped when fed into the model as a technique to improve rotational invariance and to avoid over-fitting. The arrays are passed twice through a combination of convolutional layer plus pooling layer, and then we apply a dropout layer to minimise potential over-fitting. After this, the results of both branches are flattened and concatenated, and we use a series of fully-connected dense layers until we reduce the output to N values where N is the number of mass profile points we intend to fit (in this case, 10). We again apply dropout between specific layers. Also, we concatenate the projected half-light radius ,the line-of-sight velocity dispersion of the galaxy and the 98th percentile of line-of-sight velocities to the outputs of the first dropout layer.
We utilise the outputs of the CNN to train a normalising flow model (Papamakarios et al., 2021) with the Python package ltu-ili (Ho et al., 2024a, Learning the Universe Implicit Likelihood Inference). Specifically, the 32 values from the penultimate layer of the CNN are employed as inputs to train a Masked Autoregressive Flow (Papamakarios et al., 2018, MAF) for neural posterior estimation. The model learns a series of transformations conditioned on the input data, applied to a base Gaussian distribution, to produce a posterior distribution for the target variables (dynamical enclosed masses in this study). The MAF model outputs a N-dimensional joint PDF representing the dynamical mass of a galaxy within N different radii, rather than single values for each radius. This approach offers advantages such as accounting for correlations between different dimensions of the output variables and enabling the estimation of uncertainties.
@@ 4.3 Output
The output of our CNN is a set of N values representing the galaxy’s dynamical mass enclosed within a given radius. We train the model to reproduce the galaxy masses by minimising a Mean Square Error (MSE) loss function:
Where is the -th label of the -th galaxy projection in our dataset of total size and is the neural network prediction for the same value. In our case, is the mass enclosed inside the -th radius we study for the -th galaxy projection of the dataset, numerically calculated from the galaxy simulation.
This process ensures the information encoded in the 32 neurons of the penultimate layer of the CNN model contains relevant information about the masses we aim to estimate, and therefore it is suitable to use as an input for the MAF model.
The output of the MAF is a N-dimensional joint PDF, where each dimension represents the galaxy’s dynamical mass enclosed within a given radius. The MAF also needs to undergo a training process so that it can learn the transformations that generate correct posterior distributions for the target data. In order to train a model that outputs a meaningful PDF , we use the negative log-likelihood () of the distribution as a loss function:
Where is the set of labels of the -th galaxy projection in our dataset and is the PDF generated by the MAF using transformations defined by the trainable parameters . Since the output distribution is N-dimensional, we evaluate it over a set of N labels , which are the galaxy enclosed masses within N different radii.
@@ 5.1 Mass estimators
We present the results of applying the literature mass estimators described in Section 3 to our dataset. We skip the estimator described in Peñarrubia et al. (2008) as we find the assumption of isotropic velocities to fail greatly on these simulated galaxies.
Fig. 4 shows the probability distribution function of the ratio between the enclosed mass calculated by the literature mass estimators and the true enclosed mass for the galaxies in our training dataset. We also show the results of applying literature mass estimators to galaxies from each individual suite of simulations. We find all literature mass estimators to produce reasonably precise dynamical mass estimations, showing 1- dispersion values ranging from 0.26 for Wolf and Errani through 0.30. The accuracy of estimators also varies from one to another, and the estimators acting at the innermost radii present biases towards underpredicting the enclosed mass (15% for Walker and 10% for Wolf). The trends observed are similar when analysing only AURIGA or NIHAO simulations. Nevertheless, FIRE simulations deviate slightly from the behaviour described above, exhibiting lower 1- dispersion and lower bias at inner radii mass estimations.
The varying performance of literature mass estimators on simulated galaxies depending on their properties or the specifics of the simulation is a topic worth exploring, but it is out of the scope of the present study. See Campbell et al. (2017) for such study on the Walker’s and Wolf’s estimators using APOSTLE simulations. Notably, the errors associated with each mass estimator in this work are generally larger than those reported in previous studies. This discrepancy arises from the application of these estimators to a diverse set of hydrodynamical simulations, as opposed to a dataset consisting solely of N-body simulations or mock data (e.g. Errani et al., 2018), or to a very uniform set of hydrodynamical simulations (Campbell et al., 2017; González-Samaniego et al., 2017).
@@ 5.2 CNN
We train our model separately on NIHAO and AURIGA projections of galaxies, and also on a combined dataset including both. We carefully separate our dataset in training and validation sets, making sure all projections of one singular galaxy lay only in one of the sets. We train our model with a batch size of 64, using the Adam optimiser with an initial learning rate of . We reduce the learning rate after each 10 epochs by a factor of 2, and we end the training when the validation loss has not improved for 15 epochs straight. The training typically terminates after 50 epochs. Next we extract the information from the penultimate layer of the CNN and we use it to train the MAF model created with the python package ltu-ili (Ho et al., 2024b) using 50 hidden features, 5 flow transformations, a batch size of 32 and a learning rate of .
@@
Once the model is trained, we sample the joint PDF obtained through our CNN+MAF model 1000 times for each galaxy projection. We obtain estimates of the dynamical enclosed mass at 10 different radii by calculating the median of the resulting PDF along each dimension, and we assign them uncertainties using the 16th and 84th percentiles. In Fig. 5 we show example predicted mass profiles for four galaxy projections in the validation set of the model trained on NIHAO and AURIGA galaxies together, compared to the numerical cumulative mass profile obtained by directly analysing the simulated galaxies and to existing mass estimator results. In Fig. 11 we present the full PDF obtained for each of the four galaxy projections.
@@
Just like for the literature mass estimators, we explore the accuracy of the CNN+MAF model using the ratio between the predicted and the real mass, only we now obtain predictions at several radii on the fly using the CNN+MAF model. In the left panel of Fig. 6 we show the accuracy of the model trained on NIHAO galaxies only, and we show the same for the model trained and tested on AURIGA galaxies in the right panel. We compare the model results with the accuracy of literature mass estimators in each set, and we observe that our CNN+MAF model outperforms all of them, both in terms of bias and scatter, while providing mass estimates at several radii which literature mass estimators are unable to explore. We notice that the median ratio between predicted and real enclosed mass remains constant and close to unity for both the training and validation set both for NIHAO and AURIGA galaxies. However, there are differences between the results for both simulations: our model predictions show increased scatter for NIHAO galaxies at the innermost radii, which we do not find in AURIGA. This may be a consequence of the absence of cored dark matter profiles among AURIGA galaxies, which makes dynamical mass profile shapes more uniform in comparison to NIHAO galaxies, among which we find both cusps and cores (Di Cintio et al., 2014; Tollet et al., 2016; Di Cintio et al., 2017). Also, only in the case of training on NIHAO we find more scatter on the validation set than on the training set, a hint of possible over-fitting. Constructing a mixed dataset to train on both simulations at the same time produces an intermediate result (Fig. 7), both in terms of inner scatter and possible slight over-fitting, while the model still shows no bias and improves with respect to literature mass estimators at all radii.
@@
Although all three models (trained on NIHAO only, AURIGA only, and on both at the same time) outperform literature mass estimators, not all of them generalize well across all simulated galaxies. Fig. 8 illustrates the results of applying these models to the validation sets of NIHAO and AURIGA, respectively. We find that a model trained exclusively on one simulation set struggles to make accurate predictions for the other. In contrast, the model trained on both simulation suites together is able to predict masses accurately for each individual set.
To assess whether training on both NIHAO and AURIGA galaxies at the same time improves the model’s ability to generalize to an entirely different set of simulations, we apply the three training models to FIRE galaxies and present the results in Fig. 9. The models trained solely on NIHAO and AURIGA exhibit systematic differences when applied to FIRE galaxies, particularly at the innermost radii. However, training on a combination of NIHAO and AURIGA produces more balanced predictions (although they present more scatter than literature mass estimators), indicating that expanding the training set can improve generalization. Furthermore, we observe that small changes in the train/test split lead to variations in the median predictions, indicating that the model can converge to different states. While it consistently performs well on NIHAO and AURIGA galaxies, its performance on FIRE presents some variability.
@@
As previously discussed, our model generates full posterior distributions for the enclosed masses. To evaluate the calibration of these distributions, we employ the Test of Accuracy with Random Points (Lemos et al., 2023, TARP) methodology, which assesses the statistical coverage of the posterior distributions. Statistical coverage refers to the proportion of times the true value of a parameter lies within a given credible interval of the posterior distribution. A well-calibrated posterior ensures that, for example, a 90% credible interval contains the true value approximately 90% of the time. In Fig. 10 we present the results of the TARP test for both the training and validation sets of the model trained on combined NIHAO and AURIGA galaxies. We find that the posteriors produced by our model are highly well-calibrated for galaxies in the training set, with deviations from the expected coverage being less than 3% across all confidence levels. Moreover, the calibration is nearly identical for all galaxies in the training set, regardless of their origin (NIHAO or AURIGA). For the validation set, the results are similarly robust, though deviations slightly exceed 3% for AURIGA galaxies at credibility levels around 60%. In contrast, deviations for NIHAO galaxies in the validation set, as well as for the combined NIHAO+AURIGA validation set, remain below 3%. Notably, the TARP results reveal more pronounced differences between galaxies from the two simulation suites in the validation set. Additionally, we apply the TARP test to the posteriors generated for FIRE galaxy projections. These posteriors exhibit increasing overconfidence as the credibility level rises, a trend consistent across multiple model trainings. In some cases, deviations from the expected coverage reach up to 30%. Further details and additional checks of the posterior calibration are provided in Appendix C.
@@ 6 Conclusions
In this work, we aim to evaluate the performance of existing literature galaxy mass estimators, as well as to construct a novel machine-learning model able to predict the enclosed mass of dispersion supported galaxies at any radius from line-of-sight stellar data. To do so, we use a large dataset from en extensive number of zoom-in cosmological hydrodynamical simulations from the projects NIHAO (Wang et al., 2015), AURIGA (Grand et al., 2024) and FIRE (Wetzel et al., 2023).
@@
We find existing mass estimators to work reasonably well on our dataset, producing estimates for NIHAO and AURIGA galaxies with dispersion of 0.25-0.30 on the ratio between the predicted and actual mass calculated from the simulation. We notice the two innermost tested literature mass estimators (Walker and Wolf) systematically underestimate the enclosed mass for these galaxies by more than 10%. We also note that the errors associated to literature mass estimators, when applied on our large set of hydrodynamical simulations, are larger than what quoted from previous studies, likely due to the latter using a more self-similar and uniform data set (Campbell et al., 2017; Errani et al., 2018; González-Samaniego et al., 2017).
@@
We developed a machine-learning model consisting on a convolutional neural network and a masked autoregressive flow, which outperforms literature mass estimators when applied to simulations from the same suites it has been trained on. The model not only improves mass estimates at radii where mass estimators have been proposed in the literature (both in terms of scatter and bias), but it also provides estimates of the dynamical mass enclosed at any radii, where no mass estimators exist. Our model trained on a mixed dataset of NIHAO and AURIGA simulated galaxies presents less than 1% bias in recovering the true enclosed galaxy mass at any given radius, with 1- scatter below 20% for radii over the projected half-light radius. A key advantage of our model is its ability to generate full posterior distributions for the enclosed masses. These posteriors are well-calibrated across all credibility levels, meaning that the probability intervals (e.g., 68%, 90%) reliably contain the true mass values at the expected rates. Additionally, the posteriors are well-calibrated for all marginals, ensuring that the one-dimensional projections of the posterior distributions for individual parameters (e.g., mass at a specific radius) are also statistically reliable. This calibration highlights the robustness and versatility of our model for dynamical mass estimation.
@@
We trained our model separately on galaxy projections from the NIHAO and AURIGA projects, and we find similarly good results for both of them. One key difference is the predictions of enclosed masses at inner radii show more scatter for NIHAO galaxies, which we attribute to the greater variety of inner dark matter density profile slopes in this simulation suite (Di Cintio et al., 2014; Tollet et al., 2016), compared to galaxies from the AURIGA project. We find that training exclusively on galaxies from one simulation suite and applying the model to another can lead to discrepancies in the predictions. However, when we combine NIHAO and AURIGA simulations for training and test the model on FIRE galaxies, the model still captures the overall trends despite some remaining variability. Although the scatter is higher compared to literature mass estimators, this outcome highlights opportunities for further refining the model, like the application of domain adaptation techniques.
@@
The lack of generalization power of the model might be due to different factors. Since we provide the model with the half-light radius and stellar velocity dispersion along the line-of-sight, and given that these properties can vary across different simulations (as shown in Fig. 2), it is possible that the model may learn biased relationships between these features and the enclosed mass values. Another possibility is there is a bias in the maps used as inputs for the CNN. Biases could appear as a consequence of the varying spatial resolution of simulations, and also by a sub-optimal kernel width choice for transforming the distribution of projected stellar positions and velocities into a PDF we can sample on arrays.
At this stage, our CNN-based model can be applied to real galaxies, but this relies on the assumption that our simulation training set closely matches the properties of the observed galaxies.
@@
In future works, in order to improve the model’s generalizability, we plan on creating a new model by following three different strategies. The first one consists in changing the neural network model used from a CNN to a Graph Neural Network (Scarselli et al., 2009), which eliminates the need to use a gaussian kernel for creating the inputs. The next option revolves around changing the normalizing flow model to one which predicts a likelihood (Neural Likelihood Estimation, Alsing et al., 2018, 2019) rather than directly outputting the posterior distribution. This would allow us to discard predictions on a new set of data never seen by the model if we find the likelihood value to be too low. Finally, another option is to apply domain adaptation techniques like Domain Adversarial Neural Networks (Ganin et al., 2016), which aim to force the model to use information which is invariant through different domains (simulation suites), improving its generalization ability.

### v4

@@ 4 Neural network
We used the python package pytorch (Paszke et al., 2019) to construct a CNN based on the one used in Expósito-Márquez et al. (2023).
@@ 4.1 Input
We constructed the inputs using the x and y line-of-sight projected positions of stars in the sky, and their velocities along the line of sight (), for each projection of each galaxy in the dataset. Additionally, we used the projected half-light radius (), the 98th percentile of the line-of-sight velocity (), and the line-of-sight velocity dispersion of stars (). We calculated these quantities for each projection using only the stars in the sub-sample corresponding to the projection, rather than all stars in the galaxy.
We chose to use two different 2D arrays with dimensions 256x256 as inputs. The arrays represent samples of 2D joint PDFs of projected stellar data. We constructed the PDFs using bivariate kernel density estimation (KDE), which is described in Appendix A. The input arrays were as follows:
Input 1: A 2D PDF of the projected positions of stars in the sky, normalised by (the two dimensions of the array are {x/, y/}). We sampled the PDF in a grid between -2.4 and 2.4 in both dimensions to create a 256x256 array.
Input 2: A 2D PDF in the space {/, /} (where ). We sampled the PDF in a grid between -2.4 and 2.4 in the / dimension and between -1 and 1 in the / dimension, obtaining a 256x256 array.
Input 3: Since literature mass estimators make use of line-of-sight velocity dispersion and projected half-light radius (), we decided to also input these values to the final layers of the network. On top of that, we included the value , as we used it for scaling one dimension of the array in input 2. We structured these data in a 1D array of size 3.
@@ 4.2 Architecture
The architecture of the CNN is shown in in Fig. 3. It is formed by two branches that take as inputs both arrays described in Sect. 4.1. The first input (projected position map {x/, y/}) is randomly rotated and flipped when fed into the model as a technique to improve rotational invariance and to avoid over-fitting. The arrays were passed twice through a combination of convolutional layer plus pooling layer, and then we applied a dropout layer to minimise potential over-fitting. After this, the results of both branches were flattened and concatenated, and we used a series of fully connected dense layers until we reduced the output to N values, where N is the number of mass profile points we intended to fit (in this case, 10). We again applied dropout between specific layers. Also, we concatenated the projected half-light radius (), the line-of-sight velocity dispersion of the galaxy (), and the 98th percentile of line-of-sight velocities () to the outputs of the first dropout layer.
We utilised the outputs of the CNN to train a normalising flow model (Papamakarios et al., 2021) with the Python package ltu-ili (Learning the Universe Implicit Likelihood Inference; Ho et al., 2024a). Specifically, the 32 values from the penultimate layer of the CNN were employed as inputs to train a masked autoregressive flow (MAF; Papamakarios et al., 2018) for neural posterior estimation. The model learns a series of transformations conditioned on the input data, applied to a base Gaussian distribution, to produce a posterior distribution for the target variables (dynamical enclosed masses in this study). The MAF model outputs a N-dimensional joint PDF representing the dynamical mass of a galaxy within N different radii, rather than single values for each radius. This approach offers advantages such as accounting for correlations between different dimensions of the output variables and enabling the estimation of uncertainties.
@@ 4.3 Output
The output of our CNN is a set of N values representing the galaxy’s dynamical mass enclosed within a given radius. We trained the model to reproduce the galaxy masses by minimising a mean square error (MSE) loss function:
where is the -th label of the -th galaxy projection in our dataset of total size and is the neural network prediction for the same value. In our case, is the mass enclosed inside the -th radius we studied for the -th galaxy projection of the dataset, numerically calculated from the galaxy simulation. This process ensures the information encoded in the 32 neurons of the penultimate layer of the CNN model contains relevant information about the masses we aimed to estimate, and therefore it is suitable to use as an input for the MAF model.
The output of the MAF is a N-dimensional joint PDF, where each dimension represents the galaxy’s dynamical mass enclosed within a given radius. The MAF also needs to undergo a training process so that it can learn the transformations that generate correct posterior distributions for the target data. In order to train a model that outputs a meaningful PDF (), we used the negative log-likelihood () of the distribution as a loss function:
where is the set of labels of the -th galaxy projection in our dataset and is the PDF generated by the MAF using transformations defined by the trainable parameters . Since the output distribution is N-dimensional, we evaluated it over a set of N labels , which are the galaxy enclosed masses within N different radii.
@@ 5.1 Mass estimators
We present the results of applying the literature mass estimators described in Sect. 3 to our dataset. We skipped the estimator described in Peñarrubia et al. (2008) as we find the assumption of isotropic velocities does not apply to these simulated galaxies.
Figure 4 shows the PDF of the ratio between the enclosed mass calculated by the literature mass estimators and the true enclosed mass for the galaxies in our training dataset. We also show the results of applying literature mass estimators to galaxies from each individual suite of simulations. We find all literature mass estimators to produce reasonably precise dynamical mass estimations, showing 1 dispersion values ranging from 0.26 for Wolf and Errani through 0.30. The accuracy of estimators also varies from one to another, and the estimators acting at the innermost radii present biases towards under-predicting the enclosed mass (15% for Walker and 10% for Wolf). The trends observed are similar when analysing only AURIGA or NIHAO simulations. Nevertheless, FIRE simulations deviate slightly from the behaviour described above, exhibiting lower 1 dispersion and lower bias at inner radii mass estimations.
The varying performance of literature mass estimators on simulated galaxies depending on their properties or the specifics of the simulation is a topic worth exploring, but it is out of the scope of the present study. See Campbell et al. (2017) for one such study on the Walker’s and Wolf’s estimators using APOSTLE simulations. Notably, the errors associated with each mass estimator in this work are generally larger than those reported in previous studies. This discrepancy arises from the application of these estimators to a diverse set of hydrodynamical simulations, as opposed to a dataset consisting solely of N-body simulations or mock data (e.g. Errani et al., 2018), or to a very uniform set of hydrodynamical simulations (Campbell et al., 2017; González-Samaniego et al., 2017).
@@ 5.2 CNN
We trained our model separately on NIHAO and AURIGA projections of galaxies, and also on a combined dataset. We carefully separated our dataset into training and validation sets, ensuring that all projections of a single galaxy are included in only one of the sets. We trained our model with a batch size of 128 using the Adam optimiser with an initial learning rate of . We reduced the learning rate by a factor of 2 after 10 epochs without improvement of the validation loss, and we ended the training when the validation loss had not improved for 25 epochs straight. The training typically terminated after 100 epochs. Next we extracted the information from the penultimate layer of the CNN and used it to train the MAF model created with the python package ltu-ili (Ho et al., 2024b) using 32 hidden features, 4 flow transformations, a batch size of 64, and a learning rate of .
@@
Once the model was trained, we sampled the joint PDF obtained through our CNN+MAF model 1000 times for each galaxy projection. We obtained estimates of the dynamical enclosed mass at ten different radii by calculating the median of the resulting PDF along each dimension, and we assigned them uncertainties using the 16th and 84th percentiles. In Fig. 5 we show example predicted mass profiles for four galaxy projections in the validation sets of NIHAO and AURIGA simulations, compared to the numerical cumulative mass profile obtained by directly analysing the simulated galaxies and to existing mass estimator results. In Fig. 11 we present the full PDF obtained for each of the four galaxy projections.
@@
Just like for the literature mass estimators, we explored the accuracy of the CNN+MAF model using the ratio between the predicted and the real mass, only we now obtained predictions at several radii on the fly using the CNN+MAF model. In the left panel of Fig. 6 we show the accuracy of the model trained on NIHAO galaxies only, and we show the same for the model trained and tested on AURIGA galaxies in the right panel. We compared the model results with the accuracy of literature mass estimators in each set, and we observe that our CNN+MAF model outperforms all of them, in terms of both bias and scatter, while providing mass estimates at several radii that literature mass estimators are unable to explore. We notice that the median ratio between predicted and real enclosed mass remains constant and close to unity for both the training and validation set both for NIHAO and AURIGA galaxies. However, there are differences between the results for both simulations: our model predictions show increased scatter for NIHAO galaxies at the innermost radii, which we do not find in AURIGA. This may be a consequence of the absence of cored dark matter profiles among AURIGA galaxies, which makes dynamical mass profile shapes more uniform in comparison to NIHAO galaxies, among which we find both cusps and cores (Di Cintio et al., 2014; Tollet et al., 2016; Di Cintio et al., 2017). In both models we find more scatter on the validation set than on the training set, a hint of possible over-fitting. Constructing a mixed dataset to train on both simulations at the same time produces an intermediate result (Fig. 7), both in terms of inner scatter and possible slight over-fitting, while the model still shows no bias and improves with respect to literature mass estimators at all radii.
@@
Although all three models (trained on NIHAO only, on AURIGA only, and on both at the same time) outperform literature mass estimators, not all of them generalize well across all simulated galaxies. Figure 8 illustrates the results of applying these models to the validation sets of NIHAO and AURIGA, respectively. We find that a model trained exclusively on one simulation set struggles to make accurate predictions for the other. In contrast, the model trained on both simulation suites together is able to predict masses accurately for each individual set.
To assess whether training on both NIHAO and AURIGA galaxies at the same time improves the model’s ability to generalise to an entirely different set of simulations, we applied the three training models to FIRE galaxies and present the results in Fig. 9. The model trained exclusively on AURIGA galaxies exhibits systematic discrepancies when applied to the FIRE dataset, particularly at small radii. In contrast, the model trained on NIHAO galaxies demonstrates improved performance on the FIRE testing set, but still shows a tendency to overestimate masses by 20 to 25% on average. This deviation is notably larger than that of traditional mass estimators, whose bias remains under 5% (except 7% for the Errani et al. (2018) mass estimator). Despite this bias, it achieves comparable scatter relative to traditional mass estimators. We attribute the better performance of the model trained on NIHAO galaxies to the structural similarity between NIHAO and FIRE simulations, both of which produce cored dark matter density profiles, a feature not captured by AURIGA simulations. A model trained on a combined dataset of NIHAO and AURIGA galaxies further decreases the scatter compared to the NIHAO-only model, although predictions are still similarly biased.
In Table 1 we present a summary of all three models (trained exclusively on NIHAO, exclusively on AURIGA, and jointly on both simulation suites) when evaluated independently on each simulation suite. We assessed model performance using the MSE between the predicted and true values of enclosed mass at all radial bins. As a baseline for comparison, we also report the MSE obtained by assigning, to each galaxy in the validation set, the mean enclosed mass at each radius computed over all galaxies in that validation set.
As presented in Table 1, the model trained jointly on NIHAO and AURIGA galaxies performs best on FIRE galaxies. In Fig. 12 we plot the median model prediction for 10 different bins of true values of the log10 enclosed mass at each radius. We find the model overestimates masses most notably for the least massive galaxies, which account for a high fraction of the FIRE testing set.
| Model | ||||
| Dataset | NIHAO | AURIGA | NIHAO+AURIGA | Baseline |
| NIHAO | 0.010 | 0.032 | 0.011 | 0.316 |
| AURIGA | 0.021 | 0.008 | 0.010 | 0.265 |
| FIRE | 0.016 | 0.034 | 0.015 | 0.326 |
@@
As previously discussed, our model generates full posterior distributions for the enclosed masses. To evaluate the calibration of these distributions, we employed the test of accuracy with random points (TARP) methodology (Lemos et al., 2023), which assesses the statistical coverage of the posterior distributions. Statistical coverage refers to the proportion of times the true value of a parameter lies within a given credible interval of the posterior distribution. A well-calibrated posterior ensures that, for example, a 90% credible interval contains the true value approximately 90% of the time. Figure 10 presents the results of the TARP test for both the training and validation sets of the model trained on the combined NIHAO and AURIGA datasets. The model demonstrates well-calibrated posterior distributions for galaxies in the training set, with deviations from the expected coverage remaining below 3% across all credibility levels. Calibration remains accurate when the model is evaluated separately on NIHAO and AURIGA galaxies within the training set, although for AURIGA galaxies, the 3% deviation threshold is slightly exceeded at credibility levels near 50%. The model tends to produce slightly overconfident predictions for NIHAO galaxies and more conservative estimates for AURIGA galaxies. The validation set exhibits similarly robust results, though the TARP test reveals more pronounced differences between the two simulation suites: predictions are more overconfident for AURIGA galaxies and more conservative for NIHAO galaxies.
We also applied the TARP test to the posteriors generated for FIRE galaxy projections by applying the same model. These posteriors reveal excessive overconfidence. In some cases, deviations from the expected coverage reach up to 20%.
Further details and additional checks of the posterior calibration are provided in Appendix D. In Fig. 13 we show the results of marginal TARP tests for each dimension of the posteriors, and in Fig. 14 we plot the radial profile of median posterior uncertainty.
@@ 6 Conclusions
In this work we evaluated the performance of existing literature galaxy mass estimators and constructed a novel machine-learning model able to predict the enclosed mass of dispersion-supported galaxies at any radius from line-of-sight stellar data. To do so, we used a large dataset from an extensive number of zoom-in cosmological hydrodynamical simulations from the projects NIHAO (Wang et al., 2015), AURIGA (Grand et al., 2024), and FIRE (Wetzel et al., 2023).
@@
We find that existing mass estimators work reasonably well on our dataset, producing estimates for NIHAO and AURIGA galaxies with dispersions of 0.25-0.30 on the ratio between the predicted and actual mass calculated from the simulation. The two innermost tested literature mass estimators (Walker and Wolf) systematically underestimate the enclosed mass for these galaxies by more than 10%. We also note that the errors associated with literature mass estimators, when applied to our large set of hydrodynamical simulations, are larger than what is quoted in previous studies, likely due to such studies using a more self-similar and uniform dataset (Campbell et al., 2017; Errani et al., 2018; González-Samaniego et al., 2017).
@@
We developed a machine-learning model consisting of a CNN and a MAF that outperforms literature mass estimators when applied to simulations from the same suites it has been trained on. The model not only improves mass estimates at radii where mass estimators have been proposed in the literature (in terms of both scatter and bias), but it also provides estimates of the dynamical mass enclosed at any radii, where no mass estimators exist. Our model trained on a mixed dataset of NIHAO and AURIGA simulated galaxies has a less than 1% bias in recovering the true enclosed galaxy mass at any given radius, with 1 scatter below 21% for radii over the projected half-light radius. A key advantage of our model is its ability to generate full posterior distributions for the enclosed masses. These posteriors are well calibrated across all credibility levels, meaning that the probability intervals (e.g. 68% and 90%) reliably contain the true mass values at the expected rates. Additionally, the posteriors are well calibrated for all marginals, ensuring that the 1D projections of the posterior distributions for individual parameters (e.g. mass at a specific radius) are also statistically reliable. This calibration highlights the robustness and versatility of our model for dynamical mass estimation.
@@
We trained our model separately on galaxy projections from the NIHAO and AURIGA projects, and we find similarly good results for the two. One key difference is that the predictions of enclosed masses at inner radii show more scatter for NIHAO galaxies, which we attribute to the greater variety of inner dark matter density profile slopes in this simulation suite (Di Cintio et al., 2014; Tollet et al., 2016) compared to AURIGA galaxies. We find that training exclusively on galaxies from one simulation suite and applying the model to another can lead to discrepancies in the predictions. However, when we combine NIHAO and AURIGA simulations for training and test the model on FIRE galaxies, the model captures the overall trends despite some remaining variability. Although the bias is larger compared to literature mass estimators (20% versus 5%), this outcome highlights opportunities for further refining the model, for example through the application of domain adaptation techniques.
@@
The model’s lack of generalisation power might be due to different factors. Since we provide the model with the half-light radius and stellar velocity dispersion along the line of sight, and given that these properties can vary across different simulations (as shown in Fig. 2), it is possible that the model learns biased relationships between these features and the enclosed mass values. Another possibility is that there is a bias in the maps used as inputs for the CNN. Biases could appear as a consequence of the varying spatial resolution of simulations, and also from a sub-optimal kernel width choice for transforming the distribution of projected stellar positions and velocities into a PDF we can sample on arrays.
At this stage, our CNN-based model can only be applied with caution to real galaxies due to the assumption that our simulation training set matches the key properties of the observed galaxies. Some discrepancies persist between hydrodynamical simulation suites and observations. For example, FIRE-2 dwarf galaxies tend to be more dispersion-dominated than their observed counterparts (El-Badry et al., 2018), AURIGA galaxies show a systematic offset from the Faber–Jackson relation (Barrientos Acevedo et al., 2023), and NIHAO simulations underpredict central stellar densities (Arora et al., 2023). Improving the applicability of our model will therefore require training and testing on a more diverse set of simulations that reproduce key observational features across a broad range of galaxy types. An additional challenge is the realistic treatment of observational effects such as uncertainties in stellar line-of-sight velocities or selection biases in stellar samples. While important, this topic is beyond the scope of our current work.
@@
In future works, in order to improve the model’s generalisability, we plan on creating a new model by following three different strategies. The first consists in changing the neural network model used from a CNN to a graph neural network (Scarselli et al., 2009), which eliminates the need to use a Gaussian kernel when creating the inputs. The next option revolves around changing the normalising flow model to one that predicts a likelihood (neural likelihood estimation; Alsing et al., 2018, 2019) rather than directly outputting the posterior distribution. This would allow us to discard predictions on a new set of data never seen by the model if we find the likelihood value to be too low. Finally, another option is to apply domain adaptation techniques like domain adversarial neural networks (Ganin et al., 2016), which aim to force the model to use information that is invariant through different domains (simulation suites), improving its generalisation ability.